OpenStudy (anonymous):

Simplify the given expression [(7g)^4]^6

OpenStudy (anonymous):

\[(7g ^{4})^{6}\] is this your question

OpenStudy (anonymous):

7g^24

OpenStudy (anonymous):

??

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[(x^a)^b = x^{ab}\]So, \[[(7g)^4]^6\]\[=[(7g)]^{4\times6}\]\[=(7g)^{24}\]\[=...\]

OpenStudy (anonymous):

\[\large \color{green}{(a^m)^n = a^{mn}}\]